OpenStudy (anonymous):

Factor theorem Let P(x) be a polynomial Prove that a.)If P(c)= 0 then x-c is a factor of P(x) b.) If x-c is a factor of P(x then P(c)= 0

OpenStudy (anonymous):

@ganeshie8 pls help

OpenStudy (anonymous):

@phi pls help

OpenStudy (phi):

according to http://www.purplemath.com/modules/factrthm.htm the "remainder theorem" states that you can write a polynomial p(x) as p(x) = (x – a)q(x) + r(x) If x – a is indeed a factor of p(x), then the remainder after division by x – a will be zero. That is: p(x) = (x – a)q(x) The factor theorem states that if you get a zero remainder (r(x) up above is 0) then (x-a) is a factor of p(x)

OpenStudy (phi):

for a) let p(x)= 0 factor p(x) into its factors (x-a1)(x-a2)...(x-an) so we have (x-a1)(x-a2)...(x-an) =0 the solution are the n roots x= a1, x= a2, .... x= an that means if p(c)=0 then c is one of these n roots, say c= a1 which also means (x-c) is a factor of p(x) b) If x-c is a factor of P(x) then P(c)= 0 Use the remainder theorem to write P(x)= (x-c)Q(x) + R(x) Now, by the factor theorem, if x-c is a factor of P(x), then R(x) is zero, therefor P(x)= (x-c)Q(x) if x= c then (x-c)=0 and P(c)= 0 * Q(x)= 0